# Exponential mixing of random interval diffeomorphisms

### v1

@@ Abstract.
We consider a finite number of orientation preserving $C^{2}$ interval diffeomorphisms and apply them randomly in such a way that the expected Lyapunov exponents at the boundary points are positive. We prove the exponential mixing, with respect to the unique stationary measure supported on the interior of the interval. The key step is to show the exponential synchronization in average.
@@
synchronization, mixing, interval diffeomorphisms, iterated function system
@@ 1. Introduction
This paper is devoted to systems of orientation preserving homeomorphisms $f_{1},\cdots,f_{m}$ of $M=[0,1]$ with a probability vector $(p_{1},\cdots,p_{m})$ assigned to them. We investigate the ergodic properties of this system when the functions are applied randomly in i.i.d. fashion. Such systems drew a lot of attention of mathematical community because of their dynamical properties (e.g. [DR07], [JD24]), and their relevance in the study of group representations on manifolds and (e.g. [Nav11], [Nav18]).
Throughout the paper =$\mathbb{N}=\{1,2,\ldots\}$, $\Sigma=\{1,\cdots,m\}^{\mathbb{N}}$, $\mathcal{F}$ is the product $\sigma$-field, $(\mathcal{F}_{n})\subseteq\mathcal{F}$ is the standard filtration. The measurable space $(\Sigma,\mathcal{F})$ is equipped with the infinite product of the measure $(p_{1},\cdots,p_{m})$ denoted by $\mathbb{P}$. The average with respect to that measure is denoted by $\mathbb{E}$. Let $f^{n}_{\omega}(x)=f_{\omega_{n}}\circ\cdots\circ f_{\omega_{1}}(x)$ for $n\geq 1$, $\omega=(\omega_{1},\omega_{2},\cdots)$ and $f^{0}_{\omega}(x)=x$. By $\theta$ we denote the left shift on $\Sigma$. With that notation $(f^{n}_{\omega}(x))_{n\geq 0}$ is the Markov process with evolution described above and started at $x\in M$.
@@
Those operators are obey the relation
[…]
@@
In the case when $M=\mathbb{S}^{1}$ and both actions by $f_{1},\cdots,f_{m}$ and $f^{-1}_{1},\cdots,f^{-1}_{m}$ are minimal (i.e. there is no closed invariant proper subset of $M$ for any of those actions) then (as Antonov [Ant84] showed) either the action is topologically conjugated to a system of rotations, or the action is synchronizing ($|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\to 0$ a.s. for every $x,y\in M$) or there exists $g\in$ Homeo($M$) of a finite order $p\geq 2$ commuting with all $f_{i}$’s. In the third case the system can be factorized by identifying the orbits of $g$ to obtain another system that appears synchronizing. Malicet [Mal17] proved that the rate of synchronization is exponential without any additional smoothness assumptions, i.e. he showed that there exists $q\in(0,1)$ such that for every $x\in M$ and $\mathbb{P}$-a.e. $\omega\in\Sigma$ there exists a neighbourhood $I$ of $x$ such that $|f^{n}_{\omega}(I)|\leq q^{n}$ for every $n\geq 1$ (an alternative proof has been given in recent paper [CI25]; paper [BM24] studies the properties of a general system satisfying this condition). That condition has been used to show the central limit theorem and the law of the iterated logarithm [SZ21], [SZ20].
For a general $M$ it is not known if Malicet’s exponential synchronization implies the exponential synchronization in average, i.e. if $\mathbb{E}|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|$ decays exponentially fast (see also Question 1 in [BM24]). In the case of the circle it has been proven by Le Page [LP82] if $f_{i}$’s are elements of PSL($2,\mathbb{R}$). His method has been adapted by Gorodetski and Kleptsyn for general synchronizing systems of circle $C^{1}$ diffeomorphisms [GK20]. Some interesting results about contacting systems on a general space $M$ are provided in [GG23], [GG24].
@@
When $M=[0,1]$ there are obviously two fixed points $0,1$, so that the interesting dynamics takes place in the interior of $M$. A stationary measure $\mu$ (i.e. a fixed point of $P$) with $\mu((0,1))=1$ not necessarily exists. However, if we assume that the expected Lyapunov exponents at the boundary
@@
are positive, then there exists such a measure $\mu_{\ast}$ [GH17]. There is an easy argument that $\mu_{\ast}$ is necessarily unique then (Theorem 1 [CS20]). The existence of $\mu_{\ast}$ implies Malicet’s exponential contraction (Corollary 2.13 in [Mal17]) for the action of homeomorphisms. This result has been exploited to show the central limit theorem [CS20] and the law of the iterated logarithm [CWSS20]. If $f_{i}$’s are $C^{1}$ diffeomorphisms it has been proven independently by Gharaei, Homburg (see Theorem 4.1 in [GH17]) that $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|$ decays exponentially fast a.s. for any two $x,y\in(0,1)$. Their argument is based on the Baxendale theorem [Bax89] (see also Lemma 4.1 in [GH17]) saying that the volume Lyapunov exponent
@@
is negative. The main result of the present paper says that Le Page and Gorodetski-Kleptsyn’s results can be adapted to the case of the interval $(0,1)$. The difficulty comes from the lack of compactness.
@@
Let $f_{1},\cdots,f_{m}$, $p_{1}\cdots,p_{m}$ satisfy (A1)-(A3). There exist $C\geq 1$, $a>0$, $\alpha>0$ and $q\in(0,1)$ such that for every Hölder function $\varphi$ with $|\varphi(x)-\varphi(y)|\leq L|x-y|^{\nu}$. Then for every $x\in(0,1)$
$\bigg{|}U^{n}\varphi(x)-\int_{0}^{1}\varphi\textrm{d}\mu_{\ast}\bigg{|}\leq CL\bigg{(}\frac{a}{\min\{x,1-x\}}\bigg{)}^{\alpha}q^{\nu n}\quad\textrm{for $n\geq 1$.}$
Corollary 2.
Let $f_{1},\cdots,f_{m}$, $p_{1}\cdots,p_{m}$ satisfy (A1)-(A3). For every $p\geq 1$ there exist $q\in(0,1)$, $C\geq 1$ such that for every Hölder function $\varphi$ with $|\varphi(x)-\varphi(y)|\leq L|x-y|^{\nu}$ Then
$\bigg{\|}U^{n}\varphi-\int_{0}^{1}\varphi\textrm{d}\mu_{\ast}\bigg{\|}_{p}\leq C\max\{L,\|\varphi\|_{\infty}\}q^{\nu n}\quad\textrm{for $n\geq 1$.}$
@@
Note that the constant $q$ appearing in Corollaries is in general different from the one in Theorem 1.
@@ 2.1. Basic facts
Let us recall once again the facts that are going to be used in the sequel. Under assumptions (A1)-(A3) there exists a fixed point $\mu_{\ast}$ of the operator $P$ in (1.1) with $\mu_{\ast}((0,1))=1$. Among all measures $\mu\in\mathcal{M}$ with $\mu((0,1))=1$ measure $\mu_{\ast}$ is unique. Moreover, under (A1)-(A3) the volume Lyapunov exponent (1.5) is negative (Lemma 4.1 in [GH17]) and $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\to 0$ a.s. for any two $x,y\in(0,1)$ (Theorem 4.1 in [GH17]). In the sequel we refer to the last property as the synchronization.
@@
The following result can easily inferred from Lemma 1 in [CS20]. We give the proof here for the convenience of the reader.
@@
Moreover, for every $a>0$ sufficiently small there exists $M\geq 1$ such that the class $\mathcal{P}_{M,\alpha}$ is invariant under the action of the corresponding Markov operator $P$ and every measure supported on $[a,1-a]$ belongs to this class. Consequently, the unique stationary measure $\mu_{\ast}$ belongs to $\mathcal{P}_{M,\alpha}$.
@@
Let us consider the system of inverse functions $f_{1}^{-1},\ldots,f_{m}^{-1}$ with the same probability vector. This new system has negative expected Lyapunov exponents at 00 and $1$ by (A1)-(A3), therefore we can find numbers $a_{1},\ldots,a_{m}$, $b_{1},\ldots,b_{m}$ and $\zeta>0$ with
@@
Now fix $a\in(0,1/2)$ and $M$ given above. Take any $\mu$ supported on $[a,1-a]$ and consider the sequence
@@
We shall need the following result, which is the consequence of the uniqueness of the stationary distribution. The proof is analogous to the proof of Proposition 4.1.13 in [KH95] and is omitted.
@@
[…]
@@ 2.2. Large deviations
(#$\mathbb{P}\bigg{(}\frac{\#\{i\leq n:f^{n}_{\omega}(x)\in[a,1-a]\}}{n}\leq 3/4\bigg{)}\leq Cq^{n}$
@@
(2.4) #$\frac{\#\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{n}\leq\frac{\sigma_{1}+\cdots+\sigma_{\rho_{n}}}{\tau_{1}+\sigma_{1}+\cdots+\tau_{\rho_{n}}+\sigma_{\rho_{n}}}=\frac{\#\{i\leq t_{\rho_{n}}:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{t_{\rho_{n}}},\quad\omega\in\Sigma.$
Indeed, if $n\in[t_{\rho_{n}},s_{\rho_{n}+1})$, then #$\#\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}=\sigma_{1}+\cdots+\sigma_{\rho_{n}}$ and $\tau_{1}+\sigma_{1}+\cdots+\tau_{\rho_{n}}+\sigma_{\rho_{n}}\leq n$. If $n\in[s_{\rho_{n}},t_{\rho_{n}})$, then (2.4) follows from the inequality $a/b\leq(a+c)/(b+c)$ (valid for $0\leq a<b$, $c>0$).
@@
[…]
@@
Let us consider $(X_{n})_{n\geq 1}$ be i.i.d. so that $X_{1}$ equals $\log a_{i}$ with probability $p_{i}$, $i=1,\cdots,m$. Then $S_{n}:=X_{1}+\cdots+X_{n}\to+\infty$ a.s. as $n\to\infty$ by (ii) and the strong law of large numbers. Thus $(S_{n})$ is transient, in particular $\mathbb{P}(\bigcup_{i=1}^{\infty}\{S_{i}<0\})<1$ (Theorem 8.1 in [Kal02]). In other words, there exists $\eta>0$ such that for every $A>0$ the probability that $(S_{n})$ enters $[A,+\infty)$ before returning to 00 is greater than $\eta$.
[…]
Now let us fix $a<b$ both sufficiently small and such that $e^{A}a<b$. Define $g_{i}(x)=a_{i}x$, $i=1,\ldots,m$. For each $k\geq 1$ define $Y_{k}^{-}:=0$ if $k-1$-th return to $[a,1-a]$ is from $(1-a,a)$ and $(g^{n}_{\theta^{t_{k-1}}\omega}(a))$ visits $(0,a)$ before $(ae^{A},1)$ and $Y_{k}^{-}:=r\wedge\min\{n\geq 1:g^{n}_{\theta^{t_{k-1}}\omega}(a)>ae^{A}\}$ otherwise. Observe that the distributions of $Y_{k}^{-}$ and $Y_{A,r}^{-}$ are the same, $k\geq 1$. Moreover $(Y_{k}^{-})$ are independent, bounded by $r$ and satisfy $Y_{k}\leq\tau_{k}$ provided $f^{t_{k-1}}(x)<a$. By the Cramér-Chernoff theorem (Theorem 27.3 in [Kal02])
@@
Let $H^{-}$ denote the event that at least $k/4$ numbers among $t_{1},\cdots,t_{k}$ have the property that $f^{t_{j}-1}_{\omega}(x)$ belongs to $(0,a)$. Let $H^{+}$ denote the event that at least $k/4$ numbers among $t_{1},\cdots,t_{k}$ have the property that $f^{t_{j}-1}_{\omega}(x)$ belongs to $(1-a,1)$. Then $H^{-}\cup H^{+}=\Sigma$. Clearly
@@
#$\frac{\#\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{n}\leq\frac{1}{4},\quad n\geq 1.$
@@
(2.9) ($\mathbb{P}(\{\rho_{n}<\lambda n\}\cap C)\leq\mathbb{P}(\{\rho_{n}<\lambda n)$
@@
($\mathbb{P}(\{\sigma_{1}+\cdots+\sigma_{\rho_{n}}>\rho_{n}L_{2}\}\cap\{\rho_{n}\geq\lambda n\})$

### v2

# Exponential decay of correlations for random interval diffeomorphisms

@@ Abstract.
We consider a finite number of orientation preserving $C^{2}$ interval diffeomorphisms and apply them randomly in such a way that the expected Lyapunov exponents at the boundary points are positive. We prove the exponential decay of correlations for Lipschitz observables with respect to the unique stationary measure supported on the interior of the interval. The key step is to show the exponential synchronization in average.
@@
synchronization, mixing, decay of correlations, interval diffeomorphisms, iterated function system
@@ 1. Introduction
This paper is devoted to iterated function systems (IFS for short) consisting of orientation preserving homeomorphisms $f_{1},\cdots,f_{m}$ of the interval $[0,1]$ with a probability vector $(p_{1},\cdots,p_{m})$ assigned to them. We investigate the ergodic properties of this system when the functions are applied randomly in i.i.d. fashion. Such systems drew a lot of attention of mathematical community because of their ergodic and dynamical properties (e.g. [Dolgopyat_Krikorian_2007], [Ilyashenko_2010], [Diaz_Gelfert_2012], [Gharaei_Homburg_2017], [DeWitt_Dolgopyat_2024]) and their relevance to the study of group representations on manifolds (e.g. [Navas_2011], [Navas_2018]).
Let $M$ be an arbitrary Polish space, let $f_{1},\cdots,f_{m}:M\rightarrow M$ be continuous transformations on $M$ and let $(p_{1},\cdots,p_{m})$ be a probability vector. Throughout the paper $\mathbb{N}=\{1,2,\ldots\}$, $\Sigma=\{1,\cdots,m\}^{\mathbb{N}}$, $\mathcal{F}$ is the product $\sigma$-field on $\Sigma$, $(\mathcal{F}_{n})\subseteq\mathcal{F}$ is the standard filtration. The measurable space $(\Sigma,\mathcal{F})$ is equipped with the infinite product of the measure $(p_{1},\cdots,p_{m})$ denoted by $\mathbb{P}$. The average with respect to that measure is denoted by $\mathbb{E}$. Let $f^{n}_{\omega}(x)=f_{\omega_{n}}\circ\cdots\circ f_{\omega_{1}}(x)$ for $n\geq 1$, $\omega=(\omega_{1},\omega_{2},\cdots)$ and $f^{0}_{\omega}(x)=x$. By $\theta$ we denote the left shift on $\Sigma$. Then $(f^{n}_{\omega}(x))_{n\geq 0}$ is the Markov process started at $x\in M$ with transition kernel
(1.1) $p(x,\cdot)=\sum_{i=1}^{m}p_{i}\delta_{f_{i}(x)}(\cdot).$
@@
Those operators obey the relation
(1.4) $\int_{M}\varphi dP\mu=\int_{M}U\varphi d\mu,\quad\mu\in\mathcal{M},\varphi\in\mathcal{B}(M).$
The central notion of the present paper is synchronization. We say the system of homeomorphisms of $M$ with probability vector is synchronizing if $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\to 0$ a.s. for any $x,y\in M$. Homburg has shown [Homburg_2018], Theorem 1.1., that on every compact manifold $M$ there exists a $C^{1}$-open set of systems of $C^{2}$ diffeomorphism such that the corresponding (minimal) iterated function system is synchronizing on some open and dense subset of $M$.
@@
This property turns out to be quite common if $M$ is one-dimensional. The first result in that direction has been obtained by Antonov [Antonov_1984], whose theorem is as follows. Let $M=\mathbb{S}^{1}$ and let $f_{1},\cdots,f_{m}$ be circle homeomorphisms. Let us assume that the IFS generated by $f_{1},\cdots,f_{m}$ is minimal (i.e. there is no closed invariant proper subset of $M$) and the IFS generated by $f^{-1}_{1},\cdots,f^{-1}_{m}$ is minimal too. Then exactly one of the following properties holds: the IFS is topologically conjugated to an IFS of rotations, the action is synchronizing or there exists $g\in$ Homeo($M$) of a finite order $p\geq 2$ commuting with all $f_{i}$’s. In the third case the system can be factorized by identifying the orbits of $g$ to obtain another system that appears synchronizing. Malicet [Malicet_2017] improved Antonov’s result by showing that in the second and third case of this trichotomy the rate of synchronization is exponential (without any additional smoothness assumptions on the homeomorphisms), i.e. he showed that there exists $q\in(0,1)$ such that for every $x\in\mathbb{S}^{1}$ and $\mathbb{P}$-a.e. $\omega\in\Sigma$ there exists a neighbourhood $I$ of $x$ such that $|f^{n}_{\omega}(I)|\leq q^{n}$ for every $n\geq 1$ (an alternative proof of Malicet’s result has been given in recent paper [Choi_2025]). That condition has been used to show (under additional minimality assumption) the uniqueness of stationary measures for the corresponding Markov chain (1.1) (Corollary 2.3 [Malicet_2017]), the central limit theorem and the law of the iterated logarithm [Szarek_Zdunik_2017], [Szarek_Zdunik_2020] for additive functionals of Markov chains (1.1). Later Gelfert and Salcedo [Gelfert_Salcedo_2024] established limit laws in the case of arbitrary proximal [^1] A system is proximal if for every $x,y\in M$ there exist $\omega\in\Sigma$ and $n_{k}\to\infty$ such that $d(f^{n_{k}}_{\omega}(x),f^{n_{k}}_{\omega}(y))\to 0$. systems with so called local contraction property (see [Gelfert_Salcedo_2024] for the definition). In [Gelfert_Salcedo_2023] the authors discussed various notions of contractivity of systems on $\mathbb{S}^{1}$ and relations between them and showed that proximal systems of $C^{1}$ circle diffeomorphisms that do not have a common invariant measure satisfy
(1.5) $\mathbb{E}d(f^{n}_{\omega}(x),f^{n}_{\omega}(y))^{\beta}\leq Cd(x,y)^{\beta}q^{n}$
for some $q\in(0,1)$, $\beta\in(0,1)$, $C\geq 1$ and all pairs $x,y\in\mathbb{S}^{1}$, where $d$ stands for the Euclidean metric on $\mathbb{S}^{1}$. The same result is proven (although in slightly disguised form) in Proposition 4.18 [Gorodetski_Kleptsyn_2020]. Le Page [LePage_1982] proved this result much earlier under additional assumption that $f_{i}$’s are elements of PSL($2,\mathbb{R}$). It is not known it Malicet’s result imply (1.5) without any smoothness assumptions – it would imply the affirmative answer to Question 1 in [Barrientos_Malicet_2024] at least in the case of systems on the circle.
@@
When $M=[0,1]$ there are obviously two fixed points $0,1$, so that the interesting dynamics takes place in the interior of $M$. A stationary measure $\mu$ (i.e. a fixed point of $P$) with $\mu((0,1))=1$ does not necessarily exists. However, if we assume that the expected Lyapunov exponents at the boundary
@@
are positive, then there exists such a measure $\mu_{\ast}$ [Gharaei_Homburg_2017]. There is an easy argument that $\mu_{\ast}$ is necessarily unique then (Theorem 1 [Czudek_Szarek_2020]). The existence of $\mu_{\ast}$ implies Malicet’s exponential contraction (Corollary 2.13 in [Malicet_2017]) for the action of homeomorphisms. This result has been exploited to show the central limit theorem [Czudek_Szarek_2020] and the law of the iterated logarithm [Czudek_Wojewodka_Szarek_2020]. If $f_{i}$’s are $C^{1}$ diffeomorphisms it has been proven independently by Gharaei, Homburg (see Theorem 4.1 in [Gharaei_Homburg_2017]) that $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|$ decays exponentially fast a.s. for any two $x,y\in(0,1)$. Their argument is based on the Baxendale theorem [Baxendale_1989] (see also Lemma 4.1 in [Gharaei_Homburg_2017]) saying that the volume Lyapunov exponent
@@
is negative. The main result of the present paper says that the relation similar to (1.5) holds on the open interval $(0,1)$. The difficulty comes from the lack of compactness.
@@
To show Theorem 1 we follow the strategy from Proposition 4.18 [Gorodetski_Kleptsyn_2020] and from [LePage_1982] rather than [Gelfert_Salcedo_2023]. Roughly speaking, the idea from [Gorodetski_Kleptsyn_2020] and [LePage_1982] can be repeated on an arbitrarily large compact subinterval $a,1-a]\subseteq(0,1)$, i.e. we can show that there exist $n_{0}$, $q_{1}\in(0,1)$, and $\beta\in(0,1)$ such that $\mathbb{E}|f^{n_{0}}_{\omega}(x)-f^{n_{0}}_{\omega}(y)|^{\beta}\leq q_{1}|x-y|^{\beta}$ for every $x,y\in[a,1-a]$. The issue is that [a,1a][a,1-a] in not invariant for the semigroup action and some trajectories leave it after $n_{0}$ iterations. In order to deal with the problem we show Proposition 5, which says that the distance between points is not too much expanded at the moment of the first common return to [a,1a][a,1-a].
Theorem 1 can be used to show the following result about the decay of correlations for the averaging operator on the space of Lipschitz observables.
@@
Let $f_{1},\cdots,f_{m}$, $p_{1}\cdots,p_{m}$ satisfy (A1)-(A3). For every $p\geq 1$ there exist $q\in(0,1)$, $C\geq 1$ such that
$\bigg|\int_{0}^{1}U^{n}\varphi(x)\psi(x)d\mu_{\ast}(x)-\int_{0}^{1}\varphi(x)\textrm{d}\mu_{\ast}(x)\int_{0}^{1}\psi(x)\textrm{d}\mu_{\ast}(x)\bigg|\leq C\|\varphi\|_{Lip}\|\psi\|_{Lip}q^{n}\quad\textrm{for $n\geq 1$.}$
for every Lipschitz functions $\varphi$, $\psi$, where $\|\varphi\|_{\textrm{Lip}}=\|\varphi\|_{\infty}+\textrm{Lip}(\varphi)$ and $\textrm{Lip}(\varphi)$ is the optimal Lipschitz constant of $\varphi$.
@@
The functions in Corollary 1 are assumed to be Lipschitz with respect to the standard Euclidean metric on $[0,1]$, i.e. d(x,y)=|xy|d(x,y)=|x-y|.
Remark 2.
Note that the constants $C$, $q$ appearing in Corollary 1 are in general different from the one in Theorem 1.
@@ 2.1. Basic facts
Let us recall once again the facts that are going to be used in the sequel. Under assumptions (A1)-(A3) there exists a fixed point $\mu_{\ast}$ of the operator $P$ in (1.2) with $\mu_{\ast}((0,1))=1$. Among all measures $\mu\in\mathcal{M}$ with $\mu((0,1))=1$ measure $\mu_{\ast}$ is unique. Moreover, under (A1)-(A3) the volume Lyapunov exponent (1.7) is negative (Lemma 4.1 in [Gharaei_Homburg_2017]) and $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\to 0$ a.s. for any two $x,y\in(0,1)$ (Theorem 4.1 in [Gharaei_Homburg_2017]). In the sequel we refer to the last property as the synchronization.
@@
The following result can easily inferred from Lemma 1 in [Czudek_Szarek_2020]. We give the proof here for the convenience of the reader.
@@
Moreover, for every $a>0$ sufficiently small there exists $M\geq 1$ such that the class $\mathcal{P}_{M,\alpha}$ is invariant under the action of the corresponding Markov operator $P$ and every measure supported on [a,1a][a,1-a] belongs to this class. Consequently, the unique stationary measure $\mu_{\ast}$ belongs to $\mathcal{P}_{M,\alpha}$.
@@
Let us consider the system of inverse functions $f_{1}^{-1},\ldots,f_{m}^{-1}$ with the same probability vector. This new system has negative expected Lyapunov exponents at 0 and $1$ by (A1)-(A3), therefore we can find numbers $a_{1},\ldots,a_{m}$, $b_{1},\ldots,b_{m}$ and $\zeta>0$ with
@@
Now fix $a\in(0,1/2)$ and $M$ given above. Take any $\mu$ supported on [a,1a][a,1-a] and consider the sequence
@@
We shall need the following result, which is the consequence of the uniqueness of the stationary distribution. The proof is analogous to the proof of Proposition 4.1.13 in [Hasselblatt_Katok_1995] and is omitted.
@@
converge uniformly to $\int\varphi d\mu_{\ast}$ on every compact set $K\subseteq(0,1)$.
@@ 2.2. Large deviations
$\mathbb{P}\bigg(\frac{\#\{i\leq n:f^{n}_{\omega}(x)\in[a,1-a]\}}{n}\leq 3/4\bigg)\leq Cq^{n}$
@@
(2.4) $\frac{\#\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{n}\leq\frac{\sigma_{1}+\cdots+\sigma_{\rho_{n}}}{\tau_{1}+\sigma_{1}+\cdots+\tau_{\rho_{n}}+\sigma_{\rho_{n}}}=\frac{\#\{i\leq t_{\rho_{n}}:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{t_{\rho_{n}}},\quad\omega\in\Sigma.$
Indeed, if $n\in[t_{\rho_{n}},s_{\rho_{n}+1})$, then $\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}=\sigma_{1}+\cdots+\sigma_{\rho_{n}}$ and $\tau_{1}+\sigma_{1}+\cdots+\tau_{\rho_{n}}+\sigma_{\rho_{n}}\leq n$. If $n\in[s_{\rho_{n}},t_{\rho_{n}})$, then (2.4) follows from the inequality $a/b\leq(a+c)/(b+c)$ (valid for $0\leq a<b$, $c>0$).
@@
$\mathbb{E}\big(e^{\gamma_{1}\sigma_{k}}\ |\ \mathcal{F}_{s_{k}}\big)\leq\sum_{n=1}^{\infty}e^{\gamma_{1}n}\mathbb{P}(\sigma_{k}\geq n)\leq\sum_{n=1}^{\infty}e^{\gamma_{1}n}c^{n-1}\leq e^{L_{1}}\quad\textrm{a.s.}$
@@
Let us consider $(X_{n})_{n\geq 1}$ be i.i.d. so that $X_{1}$ equals $\log a_{i}$ with probability $p_{i}$, $i=1,\cdots,m$. Then $S_{n}:=X_{1}+\cdots+X_{n}\to+\infty$ a.s. as $n\to\infty$ by (ii) and the strong law of large numbers. Thus $(S_{n})$ is transient, in particular $\mathbb{P}(\bigcup_{i=1}^{\infty}\{S_{i}<0\})<1$ (Theorem 8.1 in [Kallenberg_2002]). In other words, there exists $\eta>0$ such that for every $A>0$ the probability that $(S_{n})$ enters $A,+\infty)$ before returning to 0 is greater than $\eta$.
For $A>0$, $r>0$ let $Y^{-}_{A,r}=0$ on $\{\textrm{$(S_{n})$ visits $(-\infty,0)$ before $[A,+\infty)$ }\}$ and $Y_{A,r}=r\wedge\min\{n\geq 1:S_{n}\geq A\}$ otherwise. By the condition proved in the preceding paragraph if $A>0$, $r>0$ are large enough then $\mathbb{E}Y^{-}_{A,r}>40L_{2}$. In the symmetric way we can define $Y^{+}_{A,r}$ when $X_{1}$ takes value $\log b_{i}$ with probability $p_{i}$, $i=1,\cdots m$. Let $A$ and $r$ be so large that both $\mathbb{E}Y^{-}_{A,r}>40L_{2}$ and $\mathbb{E}Y^{+}_{A,r}>40L_{2}$.
Now let us fix $a<b$ both sufficiently small and such that $e^{A}a<b$. Define $g_{i}(x)=a_{i}x$, $i=1,\ldots,m$. For each $k\geq 1$ define $Y_{k}^{-}:=0$ if k1k-1-th return to [a,1a][a,1-a] is from (1a,a)(1-a,a) and $(g^{n}_{\theta^{t_{k-1}}\omega}(a))$ visits $(0,a)$ before $(ae^{A},1)$ and $Y_{k}^{-}:=r\wedge\min\{n\geq 1:g^{n}_{\theta^{t_{k-1}}\omega}(a)>ae^{A}\}$ otherwise. Observe that the distributions of $Y_{k}^{-}$ and $Y_{A,r}^{-}$ are the same, $k\geq 1$. Moreover $(Y_{k}^{-})$ are independent, bounded by $r$ and satisfy $Y_{k}\leq\tau_{k}$ provided $f^{t_{k-1}}(x)<a$. By the Cramér-Chernoff theorem (Theorem 27.3 in [Kallenberg_2002])
@@
Let $H^{-}$ denote the event that at least $k/4$ numbers among $t_{1},\cdots,t_{k}$ have the property that $f^{t_{j}-1}_{\omega}(x)$ belongs to $(0,a)$. Let $H^{+}$ denote the event that at least $k/4$ numbers among $t_{1},\cdots,t_{k}$ have the property that $f^{t_{j}-1}_{\omega}(x)$ belongs to (1a,1)(1-a,1). Then $H^{-}\cup H^{+}=\Sigma$. Clearly
@@
$\frac{\#\{i\leq n:f_{\omega}^{i}(x)\not\in[a,1-a]\}}{n}\leq\frac{1}{4},\quad n\geq 1.$
@@
(2.9) $\mathbb{P}(\{\rho_{n}<\lambda n\}\cap C)\leq\mathbb{P}(\{\rho_{n}<\lambda n)$
@@
$\mathbb{P}(\{\sigma_{1}+\cdots+\sigma_{\rho_{n}}>\rho_{n}L_{2}\}\cap\{\rho_{n}\geq\lambda n\})$